OpenStudy (anonymous):

Find the limit \[\lim_{n \rightarrow \infty} \frac{ 1 }{ n } \sum_{i=1}^{n} \frac{ 1 }{ 1+(\frac{ i }{ n })^2 }\]

OpenStudy (anonymous):

I got: \[\frac{ \pi }{ 4 }\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{ 1 }{ n } \sum_{i=1}^{n} \frac{ 1 }{ 1+(\frac{ i }{ n })^2 }\]

OpenStudy (experimentx):

yes you are right .. change it into integration

OpenStudy (anonymous):

Thanks :) .

OpenStudy (experimentx):

this should be equal to \[ \int_0^1 {1 \over 1 + x^2} dx\]